# Adaptive Draft-Verification for Efficient Large Language Model Decoding

### v1

@@ 7 Conclusion
ADED improves the LLM decoding process by introducing adaptability and efficiency, significantly reducing latency and computational demands. This method achieves up to a 2.5X speedup in decoding and a 20% improvement in acceptance rates, outperforming traditional techniques. Unlike existing approaches, ADED dynamically adjusts the draft distribution using a tri-gram matrix and enhances draft quality through MCTS, eliminating the need for fine-tuning. The continuous feedback loop ensures ongoing improvements in draft generation. While ADED demonstrates robust performance across various benchmarks, future work will focus on further optimizing the adaptability mechanisms and exploring its application in more diverse real-world scenarios. Additionally, addressing potential limitations in extremely large-scale deployments will be a priority.
@@
Appendix A Advantages on Computation Effiency
@@ Appendix B Broader Impacts
Appendix C Preliminary
@@
where $p$ denotes the conditional probability distribution calculated by the LLM with parameters $\theta_{\text{large}}$. To reduce these computational burdens during inference, speculative decoding is proposed Leviathan et al. [2023]. It reduces the frequency of forward passes using $\theta_{\text{large}}$ by incorporating an auxiliary language model with fewer parameters $\theta_{\text{small}}$.
@@
Monte Carlo Tree Search:  Monte Carlo Tree Search (MCTS) Coulom [2007] is a popular algorithmic in artificial intelligence, which explores potential future states from a current decision point by building a tree of possibilities, balancing exploration of new paths and exploitation of known beneficial paths.
In the context of MCTS, each node in the tree represents a possible state, and these nodes are expanded based on the results of simulated plays. A key aspect of MCTS is how it selects nodes for further exploration, which is based on the number of visits to each node, denoted as $N(s)$. Specifically, the selection process in MCTS can be interpreted as an approximate solution to a regular optimization problemGrill et al. [2020]. The node visit count $N(s)$ acts as a regularizer, guiding the algorithm towards a balance between exploring less visited, uncertain nodes and exploiting nodes that are known to yield higher rewards. The optimization problem can be formally expressed as
@@
Appendix D Configuration of ADED
For our experiments, we use the following hyperparameters to optimize the performance of ADED: We set $t$, the threshold for the elimination of tri-gram probability, at 12 to focus only on the most prevalent trigrams. The number of iterations for Monte Carlo Tree Search (MCTS), denoted $s$, is fixed at 150. The parameters in the MCTS PUTC Score function, $c_{1}$ and $c_{2}$, are 2.414 and 8.0, respectively, to effectively balance exploration and exploitation. The search depth and the length of each retrieved continuation candidate, represented by $l$, is 4; and the number of continuation candidates, $n$, is set at 24.
Appendix E Configuration of Baselines
REST
The REST baseline uses the default settings with the following specific configurations:
@@ Appendix E Configuration of Baselines
Lookahead
The Lookahead baseline uses the default settings with the following specific configurations:
@@
Appendix F Implementation Details
@@
F.1 Monte Carlo Tree Search for Text Generation
@@
1:procedure RunMCTS($self,iterations$)
2:     $rng\leftarrow\text{initialize random number generator}$
3:     for $iter\leftarrow 1$ to $iterations$ do
4:         $node\leftarrow self.root$
5:         $state\leftarrow[node.word]$
6:         $end\leftarrow\text{False}$triangleright Selection
7:         while $\text{not }node.untried\_words\text{ is empty and }node.children\text{ exist and not }end$ do
8:              $selected\_child\leftarrow node.select\_child()$
9:              $state.append(selected\_child.word)$
10:              $node\leftarrow selected\_child$
11:              if $\text{length of }state==self.sentence\_length$ then
12:                  $end\leftarrow\text{True}$
13:              end if
14:         end $\triangleright$ Expansion
15:         if $\text{not }node.untried\_words\text{ is empty and not }end$ then
16:              $untried\_words\leftarrow node.untried\_words$
@@
18:              for $word\text{ in }untried\_words$ do
19:                  $tmp\_untried\_words\leftarrow\text{get potential words from trigram matrix}$
20:                  $child\_node\leftarrow\text{new Node}(p\_word,word,tmp\_untried\_words)$
21:                  $node.children.append(child\_node)$
22:              end for
23:              $node\leftarrow node.children.last()$
24:              $state.append(node.word)$
25:         end $\triangleright$ Simulation
26:         while $\text{length of }state<self.sentence\_length$ do
27:              $last\_word\leftarrow state.last()$
28:              $next\_words\leftarrow\text{get from trigram matrix using }last\_word$
29:              if $next\_words\text{ is not empty}$ then
30:                  select next word based on probabilities from $\text{select next word based on probabilities from }next\_words$
31:                  $state.append(\text{selected word})$
32:              else
33:                  break
34:              end if
35:         end $\triangleright$ Backpropagation
36:         $score\leftarrow 1.0$
37:         for $i\leftarrow 1$ to length of $\text{length of }state-1$ do
38:              $score\leftarrow score+\text{get score from trigram matrix}$
39:         end for
40:         $current\_node\leftarrow node$
41:         while $current\_node\text{ is not None}$ do
42:              $current\_node.visits\leftarrow current\_node.visits+1$
43:              $current\_node.score\leftarrow current\_node.score+score$
44:              $current\_node\leftarrow current\_node.parent$
45:         end while
46:     end $\triangleright$ Extract and sort top sentences
47:     $sentences\leftarrow\text{extract sentences from root}$
48:     sort $\text{sort }sentences\text{ by score}$
49:     return $sentences$
50:end procedure
@@
Expansion: If the node has untried words and the end of the sentence is not reached, child nodes are created for each untried word using potential words from the trigram matrix.
Simulation: A sequence of words is generated starting from the current state until the sentence reaches the desired length or no more words can be selected from the trigram matrix.
@@
F.2 Dynamic Adjustment of tri-gram Probabilities
@@
For each trigram in the recent portion of the token sequence, the probability is increased by the increment value, up to the specified maximum probability. If the trigram is not already in the tri-gram matrix, it is added with the increment value as its initial probability.
@@
1:procedure Adjust3Gram($tokens,maxLength,increment,maxProb$)
2:     $n\leftarrow\text{length of }tokens$
3:     $startIndex\leftarrow\max(0,n-maxLength)$
4:     $subTokens\leftarrow tokens[startIndex:n]$
5:     for $i\leftarrow 2$ to length of $\text{length of }subTokens$ do
6:         $triGram\leftarrow(subTokens[i-2],subTokens[i-1],subTokens[i])$
@@
8:              $currentProb\leftarrow\text{tri-gram matrix}[triGram]$
9:              $newProb\leftarrow\min(currentProb+increment,maxProb)$
@@
13:         end if
14:     end for
15:end procedure
Appendix G Additional Experimental Results
In this section, we present additional experimental results to further illustrate the effectiveness of our proposed methods. We compare the performance of Vicuna-7B and Vicuna-13B models with and without the adaptive strategy and analyze the impact of varying Monte Carlo Tree Search (MCTS) search counts on performance.
Figure  7 shows the performance comparison on the MTBench dataset for Vicuna-7B and Vicuna-13B models with and without the adaptive strategy. The results demonstrate the advantage of using the adaptive approach. For both models, the adaptive strategy significantly reduces the average exact length of the generated sequences as the number of tokens increases, indicating more efficient and accurate text generation.
(a) Vicuna-7B
(b) Vicuna-13B
Figure 7: Adaptive Strategy Comparison on MTBench: Performance comparison of Vicuna-7B (a) and Vicuna-13B (b) models with and without the adaptive strategy on the MTBench dataset, showing the advantage of using the adaptive approach.
Figure  8 presents the results for Vicuna-7B and Vicuna-13B models on the MTBench dataset, showing the impact of varying MCTS search counts on performance. For both models, increasing the number of MCTS search counts leads to improved performance, with the optimal counts varying by model size. The average exact length and latency are plotted against the number of MCTS search counts, illustrating the trade-off between performance and computational cost. As shown in the plots, there is a notable improvement in the average exact length as the search counts increase, while latency also increases, indicating a balance between the depth of search and the time taken for generation.
(a) Vicuna-7B
(b) Vicuna-13B
Figure 8: Comparison of MCTS Search Counts and Performance on MTBench: Results for Vicuna-7B (a) and Vicuna-13B (b) models on the MTBench dataset. Increasing MCTS search counts improves performance, with optimal counts varying by model size.

### v2

@@ Conclusion
ADED improves the LLM decoding process by introducing adaptability and efficiency, significantly reducing latency and computational demands. This method achieves up to a 2.5X speedup in decoding and a 20% improvement in acceptance rates, outperforming traditional techniques. Unlike existing approaches, ADED dynamically adjusts the draft distribution using a tri-gram matrix and enhances draft quality through MCTS, eliminating the need for fine-tuning. The continuous feedback loop ensures ongoing improvements in draft generation. While ADED demonstrates robust performance across various benchmarks, future work will focus on exploring its application in more diverse real-world scenarios. Additionally, addressing potential limitations in extremely large-scale deployments will be a priority.
@@
Appendix A Advantages on Computation Efficiency
@@ Appendix B Broader Impacts
Appendix C Limitations
The ADED methodology, while effective in accelerating LLM decoding, presents certain limitations primarily concerning accuracy and efficiency. Firstly, the retrieval-based approach integral to ADED, although less resource-intensive, does not yet match the accuracy levels of traditional draft model methods. This discrepancy highlights a potential trade-off between resource efficiency and optimal accuracy in language model decoding. Secondly, the use of Monte Carlo Tree Search (MCTS) in our adaptive draft-verification process, though instrumental in improving decoding speed, still incurs additional computation. As a result, while ADED achieves notable speed enhancements, the acceleration ratio remains suboptimal when compared to the theoretical maximum acceptance length. Future improvements will aim to refine these aspects to better balance speed and accuracy, potentially through more advanced optimization of the MCTS algorithm or alternative adaptive strategies.
Appendix D Ablation Study
Effect of Draft Search Score
In this section, we present additional experiments to elucidate our choice and the implications of the hyperparameters $c_{1}$ and $c_{2}$ in the Draft Search Score formula.
Our analysis is grounded in Equation 16, which demonstrates that the exploration-exploitation balance is modulated by $c_{1}$ and $c_{2}$. Specifically, $c_{1}$ positively correlates with the exploration constant $E$, while $c_{2}$ inversely affects it.
(a) Impact of $c_{1}$ on Acceptance Length
(b) Impact of $c_{2}$ on Acceptance Length
Figure 7: Relationship between Average Acceptance Length and Exploration Parameters.
Figure 7 illustrates the relationship between the average acceptance length and both $c_{1}$ and the reciprocal of $c_{2}$ ($1/c_{2}$). We observe an initial increase in performance with moderate levels of exploration, followed by a decline as the parameters extend to either extreme. This pattern suggests that both insufficient and excessive exploration can detrimentally affect model performance. This pattern indicates that both insufficient and excessive exploration levels impact the algorithm’s performance, highlighting the importance of careful adjustment of $c_{1}$ and $c_{2}$.
Summary
In Experiment section, we conduct ablation studies to explore the impact of various configurations and optimizations in ADED.
The adaptive strategy shows significant benefits across different models, notably improving accept lengths by dynamically adjusting probability distributions based on tri-gram frequencies.
Our examination of the effects of corpus size indicates that larger datasets contribute to improved accuracy, enhancing the coherence and quality of the output. However, even with smaller corpus sizes, our system demonstrates a notable increase in processing speed, underscoring its efficiency and capability to deliver accelerated performance without compromising on resource constraints.
The effectiveness of the Monte Carlo Tree Search (MCTS) is analyzed through varying search counts, demonstrating that increased iterations enhance performance while balancing computational costs.
Additionally, comparisons between different n-gram models indicate that tri-grams achieve an optimal balance of randomness and contextual relevance, significantly outperforming other configurations in terms of decoding performance.
These studies collectively affirm the critical design choices made in ADED, validating its robustness and efficiency in language model decoding.
Appendix E Preliminary
@@
where $p$ denotes the conditional probability distribution calculated by the LLM with parameters $\theta_{\text{large}}$. To reduce these computational burdens during inference, speculative decoding is proposed (Leviathan, Kalman, and Matias 2023). It reduces the frequency of forward passes using $\theta_{\text{large}}$ by incorporating an auxiliary language model with fewer parameters $\theta_{\text{small}}$.
@@
Monte Carlo Tree Search:  Monte Carlo Tree Search (MCTS(Coulom 2007) is a popular algorithmic in artificial intelligence, which explores potential future states from a current decision point by building a tree of possibilities, balancing exploration of new paths and exploitation of known beneficial paths.
In the context of MCTS, each node in the tree represents a possible state, and these nodes are expanded based on the results of simulated plays. A key aspect of MCTS is how it selects nodes for further exploration, which is based on the number of visits to each node, denoted as $N(s)$. Specifically, the selection process in MCTS can be interpreted as an approximate solution to a regular optimization problem(Grill et al. 2020). The node visit count $N(s)$ acts as a regularizer, guiding the algorithm towards a balance between exploring less visited, uncertain nodes and exploiting nodes that are known to yield higher rewards. The optimization problem can be formally expressed as
@@
PUCT Score:  The Probabilistic Upper Confidence Bound applied to Trees (PUCT) score is an adaptation of the Upper Confidence Bound (UCB) used in Monte Carlo Tree Search to optimize the exploration-exploitation trade-off in decision-making environments. The UCB formula is given by:
$\text{UCB}=Q(s,a)+\sqrt{\frac{2\log N}{N(s,a)}},$ (13)
where $Q(s,a)$ represents the average reward of taking action $a$ in state $s$, $N$ is the total number of trials, and $N(s,a)$ is the number of trials where action $a$ was taken from state $s$.
To enhance the decision-making in tree-based strategies, the UCB was extended to Upper Confidence Bound applied to Trees (UCT), which adjusts the exploration parameter dynamically:
$\text{UCT}=Q(s,a)+C\cdot\sqrt{\frac{\log N}{N(s,a)}},$ (14)
where $C$ is a variable exploration coefficient, adjusting dynamically based on the tree’s exploration needs.
The PUCT score further incorporates probabilistic assessments into the UCT framework, allowing for decisions that consider the variance in outcome distributions:
$\text{PUCT}=Q(s,a)+c\cdot P(s,a)\cdot\sqrt{\frac{\log N(s)}{1+N(s,a)}},$ (15)
where $c$ is a constant scaling the exploration term, and $P(s,a)$ represents the prior probability of selecting action $a$ in state $s$, enhancing the traditional exploration term by accounting for underlying probability distributions.
In our implementation, we use a variant of the original PUCT formula that allows for finer granularity in adjustments:
$E=C_{1}+\log\left(\frac{\sum_{b}N(s,b)+C_{2}+1}{C_{2}}\right),$ (16)
$\text{PUCT}(s,a)=Q(s,a)+E\cdot P(s,a)\cdot\frac{\sqrt{\sum_{b}N(s,b)}}{1+N(s,a)}.$ (17)
This adaptation allows our model to dynamically calibrate the exploration-exploitation balance more effectively, particularly in environments with large and complex decision spaces such as those involved in LLM decoding. The parameters $c_{1}$ and $c_{2}$ are introduced to further refine the balance between exploration and exploitation, where $c_{1}$ scales the exploration bonus and $c_{2}$ stabilizes the logarithmic term for numerical stability. This nuanced approach enhances the adaptability and efficiency of our MCTS framework, providing a substantial improvement over standard UCB and UCT methods.
Appendix F Detailed Experimental Settings
Configuration of ADED
For our experiments, we use the following hyperparameters to optimize the performance of ADED: We set $t$, the threshold for the elimination of tri-gram probability, at 12 to focus only on the most prevalent tri-grams. The number of iterations for Monte Carlo Tree Search (MCTS), denoted $s$, is fixed at 150. The parameters in the MCTS PUCT Score function, $c_{1}$ and $c_{2}$, are 32.0 and 8.0, respectively, to effectively balance exploration and exploitation. The search depth and the length of each retrieved continuation candidate, represented by $l$, is 4; and the number of continuation candidates, $n$, is set at 24.
Experimental Settings
  • For all experiments listed in Table 1, ADED uses UltraChat as the corpus, with the same configuration as mentioned in the previous section. All data presented in the table are the medians of three repeated runs.
  • For Figure 4a, ADED uses ShareGPT as the corpus with $c_{1}=2.414$.
  • For Figure 4b, ADED employs ShareGPT as the corpus.
  • For Figure 5, ADED utilizes ShareGPT as the corpus with $c_{1}=2.414$, and experiments are repeated seven times.
  • For Table 2, ADED is tested using two different corpora: ShareGPT (corpus size 121k) and UltraChat (corpus size 774k), with results reported as the median of three repetitions.
  • For Figure 5, ADED employs ShareGPT as the corpus, setting $c_{1}=2.414$.
Appendix G Configuration of Baselines
REST The REST baseline uses the default settings with the following specific configurations:
@@ Appendix G Configuration of Baselines
REST Single Thread The REST Single Thread baseline uses the same settings as REST except the Number of threads:
  • Number of threads: 1
  • Draft choice: 64
  • Datasets: UltraChat and The Stack
  • Token spans: [16, 15, 14, 13, 12, 11, 10, 9, 8, 7, 6, 5, 4, 3, 2]
Lookahead The Lookahead baseline uses the default settings with the following specific configurations:
@@
Appendix H Implementation Details
@@
Monte Carlo Tree Search for Text Generation
@@
Input: $iterations$ — Number of iterations for MCTS, $self$ — The object containing the tree structure and parameters
Output: $sentences$ — Generated sentences ranked by their scores
1:  $rng\leftarrow\text{initialize random number generator}$
2:  for $iter\leftarrow 1$ to $iterations$ do
3:     $node\leftarrow self.root$
4:     $state\leftarrow[node.word]$
5:     $end\leftarrow\text{False}$Selection}
6:     while $\text{not }node.untried\_words\text{ is empty}$ and $node.children\text{ exist}$ and not $\text{not }end$ do
7:        $selected\_child\leftarrow node.select\_child()$
8:        $state.append(selected\_child.word)$
9:        $node\leftarrow selected\_child$
10:        if $\text{length of }state==self.sentence\_length$ then
11:           $end\leftarrow\text{True}$
12:        end if
13:     end while{Expansion}
14:     if $\text{not }node.untried\_words\text{ is empty and not }end$ then
15:        $untried\_words\leftarrow node.untried\_words$
@@
17:        for $word\text{ in }untried\_words$ do
[…]
19:           $child\_node\leftarrow\text{new Node}(p\_word,word,tmp\_untried\_words)$
20:           $node.children.append(child\_node)$
21:        end for
22:        $node\leftarrow node.children.last()$
23:        $state.append(node.word)$
24:     end if{Simulation}
25:     while $\text{length of }state<self.sentence\_length$ do
26:        $last\_word\leftarrow state.last()$
27:        $next\_words\leftarrow\text{get from tri-gram matrix using }last\_word$
28:        if $next\_words\text{ is not empty}$ then
29:           select next word based on probabilities from $next\_words$
30:           $state.append(\text{selected word})$
31:        else
32:           break
33:        end if
34:     end while{Backpropagation}
35:     $score\leftarrow 1.0$
36:     for $i\leftarrow 1$ to length of $\text{length of }state-1$ do
[…]
38:     end for
39:     $current\_node\leftarrow node$
40:     while $current\_node\text{ is not None}$ do
41:        $current\_node.visits\leftarrow current\_node.visits+1$
42:        $current\_node.score\leftarrow current\_node.score+score$
43:        $current\_node\leftarrow current\_node.parent$
44:     end while
45:  end for{Extract and sort top sentences}
46:  $sentences\leftarrow\text{extract sentences from root}$
47:  sort $\text{sort }sentences\text{ by score}$
48:  return $sentences$
@@
Expansion: If the node has untried words and the end of the sentence is not reached, child nodes are created for each untried word using potential words from the tri-gram matrix.
Simulation: A sequence of words is generated starting from the current state until the sentence reaches the desired length or no more words can be selected from the tri-gram matrix.
@@
Dynamic Adjustment of tri-gram Probabilities
@@
For each tri-gram in the recent portion of the token sequence, the probability is increased by the increment value, up to the specified maximum probability. If the tri-gram is not already in the tri-gram matrix, it is added with the increment value as its initial probability.
@@
Input: $tokens$ — Sequence of tokens, $maxLength$ — Maximum length of tokens to consider
Parameter: $increment$ — Value by which the probability should be increased, $maxProb$ — Maximum allowable probability
Output: Updated tri-gram matrix with adjusted probabilities
1:  $n\leftarrow\text{length of }tokens$
2:  $startIndex\leftarrow\max(0,n-maxLength)$
3:  $subTokens\leftarrow tokens[startIndex:n]$
4:  for $i\leftarrow 2$ to length of $\text{length of }subTokens$ do
5:     $triGram\leftarrow(subTokens[i-2],subTokens[i-1],subTokens[i])$
@@
7:        $currentProb\leftarrow\text{tri-gram matrix}[triGram]$
8:        $newProb\leftarrow\min(currentProb+increment,maxProb)$
@@
12:     end if
13:  end for
14:  return Updated tri-gram matrix
Appendix I Detailed Explanation of Core Concepts and Interactions
This section provides a detailed explanation of the relationships and interactions among the core components of our methodology: the Large Language Model (LLM), the LLM representative tri-gram matrix, Monte Carlo Tree Search (MCTS), and the final output decoding. These components are integral to our Adaptive Draft-Verification for Efficient LLM Decoding (ADED) framework.
LLM Representative Tri-gram Matrix
The LLM representative tri-gram matrix is a crucial component of our ADED framework. It serves as a condensed representation of the LLM’s knowledge, capturing the conditional probabilities of token sequences based on their occurrence within the training corpus. This matrix is dynamically updated during the decoding process to reflect the changing context and to better predict subsequent tokens.
Monte Carlo Tree Search (MCTS)
MCTS is employed to explore potential draft outputs based on the probabilities indicated by the tri-gram matrix. It balances exploration of new, potentially high-reward token sequences with the exploitation of known, high-probability paths. This ensures a diverse yet accurate set of drafts that are likely to align closely with the true output distribution of the LLM.
Interaction Between the Tri-gram Matrix and MCTS
The tri-gram matrix provides the initial probability estimates that guide the MCTS in selecting the most promising token sequences during the draft generation phase. As MCTS explores different paths, it accumulates data about the success of various sequences, which in turn feedbacks into the tri-gram matrix. This feedback loop allows the tri-gram matrix to evolve and adapt, improving its accuracy and the efficiency of the draft verification process.
Final Output Decoding
Once the drafts have been generated and evaluated using MCTS, the final output decoding phase begins. This process involves comparing the drafts against the output distribution of the LLM using tree attention to select the sequence that best matches the expected outcomes. The selected draft then becomes the final output of the model for the given input context.
Summary
To aid in understanding, below is a conceptual diagram illustrating the interactions among these components. The following Figure  8 shows the flow of data and decisions from the initial input through the tri-gram matrix, MCTS, and finally to the output.
Figure 8: Data Flow and Interaction Among Core Algorithm Components. This figure details the sequential data processing steps within our ADED. Initially, the input tokens are processed to compute their tri-grams, which are then utilized to update the tri-gram matrix. Subsequently, the updated tri-gram matrix, in conjunction with the last two tokens of the input, guides the Monte Carlo Tree Search (MCTS) in retrieving potential token sequences. The retrieval results are then ranked, and the top-k sequences are selected and appended to the original input. These extended sequences are fed into the Large Language Model (LLM) for further processing.
Figure 9: Validation of Retrieval Results in Large Language Model.
Additionally, Figure 9 complements this explanation, depicting the validation process used to assess the accuracy of retrieval results using tree attention.